OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13) 4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5

OpenStudy (anonymous):

substitute the points and see which eq satisfies!

OpenStudy (anonymous):

4x +y =5

OpenStudy (anonymous):

y2 - y1 over x2-x1 = -4 , serves as your slope y-y1=m(x-x1) you can perform that

OpenStudy (anonymous):

@Aaron0620 did you understand it

OpenStudy (anonymous):

yes, very much. thank you

OpenStudy (anonymous):

ur welcome. what year are you now?